A square sheat of steal has a side of 15cm at 00 C. Determine its area at 400C. (Linear expansivity of steel = 1.1x10-5 0C-1
1
Expert's answer
2020-06-03T12:14:47-0400

Solution.

"L_0=15cm=0.15m;"

"T_0=0^oC;"

"T_1=400^oC;"

"\alpha_L=1.1\sdot10^{-5 }\dfrac{1}{oC}";

"\alpha_L=\dfrac{1}{L}\dfrac{\Delta L}{\Delta T};\implies" "\Delta L=\alpha_L\Delta T L;"

"\Delta L=1.1\sdot10^{-5}\dfrac{1}{oC}\sdot400^oC\sdot0.15m=66\sdot10^{-5}m;"

"L=L_0+\Delta L;"

"L=0.15m+0.00066m=0.15066m;"

"S=L^2;"

"S=0.15066m\sdot0.15066m=0.0227m^2;"

Answer:"S=0.0227m^2."